Can you please help me solve this problem: A manufacturer sells a product at $8 per unit, selling all that is produced. Fixed cost is $5000 and variable cost per unit is 22/9 dollars. Find the total output and revenue at the break-even point. Find the ouput required to obtain a profit of $10,000.
Thank you, I really appreicate your help

Answer by checkley71(8403)

8X=5000+22/9X
8X-22/9X=5000
(72X-22X)/9=5000
50X/9=5000 CROSS MULTIPLY
50X=50000*9
50X=45000
X=45000/50
X=900 NUMBER OF UNITS NEEDED TO BE BUILT & SOLD TO BREAK EVEN
8X-22/9X=10000
50X/9=10000
50X=90000
X=9000/50
X=1800 ADDITIONAL UNITS NEED TO BE BUILT & SOLD FOR A PROFIT OF $10,000.
THUS THE TOTAL BUILT & SOLD WOULD BE 900+1800=2700 UNITS.